What are the steps to solving equations by multipling by a reciprical and by a common denominator?
Answer by apachemexican1(64)

THIS IS VERY SIMPLE
i will show u an example
a/2=4
first be sure to add the same number to both sides
2/1*a/2=4*2/1
now you would cancel out the 2's
you would be left with
a/1=8/1
now you know that when you have any number or letter over 1 the answer is that number or letter
so the answer is
a=8
good luck